40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5474     Entry time: Tue Sep 20 03:02:23 2011
Author: Keiko 
Type: Update 
Category: LSC 
Subject: locking activity tonight 

 Keiko, Anamaria, Koji

We were not able to establish the stable DRMI tonight. We could lock MICH and PRCL quite OK, and lock the three degrees of freedom at somewhere strange for several seconds quite easily, but the proper DRMI lock was not obtained.

When MICH and PRC are locked to the carrier, REFL DC PD reading dropps from ~3000 counts to 2600~2700 counts as REFL beam is absorbed to PRC. We'll try to lock PRC to sidebands - but flipping gain sign didn't work today, although it worked a few days ago. 

POP beam (monitor) is useful to align PRM.

ELOG V3.1.3-